DIRC Operations Weekly Report
Period: Wednesday 2007/6/27 - Wednesday 2007/7/4
Oncall: Justine Serrano
Fast Monitoring: Justine Serrano, Jose Benitez
Luminosities this week
PEP peak lumi: ~1 x 1034
(last week: ~1 x 1034)
PEP delivery: 3.29 fb-1
(last weeks: 1.92 fb-1)
BaBar integrated lumi:3.17 fb-1
(last weeks: 1.76 fb-1)
Run 6 total delivery: 54.31 fb-1
Run 6 total BaBar recorded: 52.22 fb-1 total, including 4.18 fb-1 off-peak data
PEP activities/issues this week
06/27: ROD. PEP people remove two flexible flanges in arc5. Beams back in the night.
BaBar activities/issues this week
06/27: ROD. Two FEAs from the DCH has been changed succesfully.
DIRC activities/issues this week
06/27: ROD Activities:
- X-Trees in sector 0 and sector 7 unplugged.
- Replaced DFB #1 in sector 6.
- Replaced DCC in sector 5.
- Before the ROD, one calibration failed. This was due to a bad board in the master timing crate, Chris O'Grady replaced it in the afternoon.
- We took calibrations and cosmics, everythings looked OK. Nicolas unmasked the alarms corresponding to the sector 5 monitoring in epics and in the alarm handler and made the DIRC runnable again by unbypassing sector 5 monitoring.
See DIRC logbook for details
Lumi and Background this week
- Unmasked 8 channels in sector 6.
- Update gains on new DFB in sector 6.
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
- Replace air filters on Wienner crates.
Justine and Jose rotating backup & on-call commissioning duties.